OpenStudy (annah.bee):

According to this equation, F=ma, how much force is needed to accelerate a 65-kg runner at 8m/s2? A. 520 N B. 520 m/s C. 8 m/s D. 8 N

OpenStudy (soprano.h.d0816):

Which do you think it is?

OpenStudy (annah.bee):

idk i think its A or B @soprano.h.d0816

OpenStudy (annah.bee):

@jabez177 do you know?

OpenStudy (retireed):

F=ma = Mass times Acceleration = (kg)( m/s^2) = Newtons 65 kg is mass times 8 m/s^2 is acceleration = ______ N 65-kg runner at 8m/s2

OpenStudy (annah.bee):

so are yousaying i need to multiply 65 and 8?

OpenStudy (soprano.h.d0816):

That's what I was getting from that.

OpenStudy (annah.bee):

so itsA!

OpenStudy (retireed):

If you want to find Force (in Newtons) multiply Mass by Acceleration.

OpenStudy (retireed):

Very good!

OpenStudy (annah.bee):

thank you so much :))

jabez177 (jabez177):

Yes, the answer when you want to find how much force, your answer will always be in Newtons. :)

OpenStudy (soprano.h.d0816):

Good job!!